# Priority aware grouping-based multihop routing scheme for RIS-assisted wireless networks

### v1

@@ III-C1 When $c=0$
This denotes a channel-aware delay unconstrained scenario, where $U_{i}$ always chooses the candidate with best channel condition, irrespective of its priority status. In case the connection does not involve any RIS, adaptive modulation (discussed in Section III-D) is employed and the corresponding data rate is determined.
@@ III-F Illustrative example for proposed strategy
The entire proposed strategy is presented in Fig. 2 in the form of a concise flowchart, which is illustrated below with the help of an example. As demonstrated in Fig. 1, $N$ sources $S_{1},\cdots,S_{N}$ want to communicate with their respective destination $D_{1},\cdots,D_{N}$ via IU $U_{1}$. The sources are prioritized based on their $\beta$ values, i.e., acceptable delay limit and channel condition. We consider in this work that if $U_{1}$ is idle, it agrees to act as a relay, with the higher priority users being served first. In this figure, we assume $S_{1}$ to have the highest priority and hence, it is served prior to $S_{2},\cdots,S_{N}$. Since there is no direct $U_{1}\rightarrow U_{2}$ or $U_{1}\rightarrow U_{3}$ LoS link, $U_{1}$ sends the packets to $U_{2}$ or $U_{3}$ via the RIS $R_{1,1}$. Here, we consider that the concerned RIS $R_{1}$ is already partitioned into four non-overlapping surfaces, $R_{1,1},R_{1,2},R_{1,3}$ and $R_{1,4}$, respectively. For communication purposes, $U_{1}$ selects $R_{1,1}$ to connect to $D_{1}$ via $U_{2}$ or $U_{3}$. It is interesting to note, that although LRD from $U_{3}$ to $D_{1}$ is greater than that from $U_{2}$ to $D_{1}$, $U_{3}$ is selected as the next hop. The reason for is attributed to the higher effective throughput in the case of $U_{1}\rightarrow R_{1}\rightarrow U_{3}$ link as compared to $U_{1}\rightarrow R_{1}\rightarrow U_{2}$. Finally, $U_{3}$ forwards the information packets to $D_{1}$, which completes the process.
@@ IV-B Delay bound for complete information transfer for a particular $S-D$ pair
Finally, Remark 4 holds for the scenario, where IUs are available for each hop and the RISs in the surroundings are not being used at all. For the scenario where the RISs are being used due to IU unavailability, the only difference will be that the service rate $\mu_{j}$ in (35) will get replaced by $\hat{R}_{j,j+1}^{r,b}$ from (16).
@@ V-A Impact of grouping and spatial correlation
By accounting for the spatial correlation, here we demonstrate the advantage of grouping in terms of the achievable data rate. A particular RIS is sub-divided into groups with identical number of patches and patch spacing being $\lambda/8$ for the correlated scenario. Moreover, only one of these groups is used for the purpose of information transfer and the results corresponding to the independent channel scenario is obtained by setting the patch spacing to $\lambda/2$. Fig. 3 illustrates the achievable data rate versus the number of groups into which the RIS has been sub-divided and this depicts the importance of incorporating the aspect of spatial correlation. Intuitively, this gain in performance can be understood from (III-A), i.e., grouping essentially implies combination of channels corresponding to the adjacent patches. Hence, the combination is more likely to be constructive when the channels are correlated. We also observe, that the ideal group size is somewhere in between the two extreme scenarios of considering the group size to be unity or the total number of patches in the RIS. This further corroborates the trade-off investigated in [18] between the channel estimation overhead and the power gain offered by the group.
Figure 3: Impact of the number of groups on achievable data rate.
For a fixed number of patches in a group, Fig. 4 depicts the effect of the patch spacing on achievable data rate. In this context, we define the quantities $dx$ and $dy$, which denote the horizontal and vertical distance between consecutive patches, respectively. We observe that for a fixed $dx$, the achievable rate increases with decreasing $dy$ and vice-versa; for example, note the performance gap at $dx=\lambda/12$ between $dy=\lambda/12,\lambda/10,$ and $\lambda/8$. This justifies the effect of spatial correlation in (III-A), which implies that the correlation is inversely proportional to the patch spacing. On the contrary, note that this aspect of patch distance does not have any impact in the independent scenario, i.e., when $dx=dy=\lambda/2$.
Figure 4: Impact of patch spacing on achievable data rate.
@@
Figure 5: Impact of number of patches in a group on achievable data rate.
@@
Fig. 5 shows the achievable data rate, defined in (16), as a function of the number of patches in a group. In this context, we consider groups with both optimal as well as random phases, where the optimal phase shift is calculated as in (13). Moreover, as proposed in Section III-B, we also take into account the scenario when the optimal phase shift is not a continuous quantity but it is chosen from a finite set of predetermined values. Specifically, in our proposed approach, we consider the group to be $2$ and $4$ bit coded, i.e., we choose the phase shift that is closest to the optimal from a set of $4(=2^{2})$ and $16(=2^{4})$ predetermined values, respectively. The figure demonstrates that, irrespective of the scenario, the achievable rate increases with the number of patches while the performance is worst in the case of random phases. Moreover, with an increase in the number of coded bits, the performance approaches the optimal. As can be seen, the achievable data rate with $4$ coded bits almost merges with its optimal counterpart.
@@ V-B1 Data throughput $\rm D_{T}$
Figure 7: Data throughput comparison.
@@ V-B3 Energy Efficiency $E_{\rm eff}$
Fig. 7 depicts an overall increasing trend of $\rm D_{T}$ with number of RIS elements in a group, irrespective of the scheme considered. The reason behind this observation can be attributed to the fact that for hops involving a RIS, the achievable throughput is a function of the number of reflecting elements that are being used. As the proposed strategy is IU traffic aware, involves channel-dependent transmission techniques, and also considers the aspect of spatial correlation among the reflecting patches, it performs significantly better compared to its competitors. Moreover, the figure illustrates the fact that, as we are dealing with a delay-constrained scenario, an LRD-based approach is not always the best solution. It may result in a lesser number of hops (as observed in Fig. 6) but at the cost of reduced data rate due to factors such as wireless channel fluctuations. Furthermore, here also we observe the impact of having more number of coded bits in the RIS.
Figure 8: Energy consumption comparison.
@@
Fig. 8 investigates the aspect of energy consumption among the proposed strategy and the existing benchmarks. It is interesting to note that ${\rm E_{C}}$, in the context of both our strategy and the LRD-based approach, decreases with an increasing number of elements in a RIS group. The reason behind this counter-intuitive observation is that, in both these scenarios, irrespective of the group size, a single phase shift is required for the entire group and the LRD becomes a criteria only when the next hop choice is concerned. Moreover, as observed from (8), the achievable data rate increases with the group size. Hence, when we combine both these factors, it leads to this interesting insight, which can also be noted from (V-B2). On the contrary, both for MRIRS and RAND, as individual phase shifts are required at all the patches, the energy consumption increases linearly with increasing group size.
@@
Finally, Fig. 9 shows the impact of the proposed strategy on the system energy efficiency. We observe that, while ${\rm E_{eff}}$ for MRIRS and RAND monotonically decreases with the number of reflecting elements in a group, it follows an exact opposite trend in the context of our proposed strategy. The reason behind this observation is identical to the one as already described above in the discussion related to Fig. 8.
Figure 9: Energy efficiency comparison.
V-C Impact of IU Mobility
@@
Here a mobile scenario is considered, i.e., the IUs have a particular velocity in a certain direction. It is to be noted that this will have a significant effect on the performance of the proposed scheme, as the inter-IU distance changes during the process of information transfer. In this context, Fig. 10 illustrates the impact of IU mobility on the proposed framework, with the coverage being taken as $60$ m and the mobility aspect being modeled by the standard random waypoint (RWP) model [35]. Specifically, for this figure, we consider a pedestrian scenario [36], where we fix the maximum possible velocity $V_{\max}$ of the IUs, consider a random velocity chosen uniformly in $[0,V_{\max}]$ and observe its impact on the system data throughput. The figure illustrates that the performance deteriorates monotonically with increasing $V_{\max}$, which is intuitive. Such an observation is due to the mobility phenomenon, where a particular IU can move outside the coverage of another IU even during the communication process. This inevitably leads to outage, resulting in lesser data rate. Moreover, note that this figure with $V_{\max}=0$ is a special case corresponding to the static scenario, i.e., Fig. 7. Furthermore, for a fixed $V_{\max}$, ${\rm D_{T}}$ decreases with the number of coded bits, which is also inline with the observation made in Fig. 7 as well.
Figure 10: Effect of IU mobility on performance.
@@ VI Conclusion
In this paper, we proposed a novel priority-based channel aware scheduling strategy for RIS-assisted multihop D2D wireless communication, which takes into account the aspect of spatial correlation and element grouping at the RIS. The proposed strategy avoids resource wastage by exploiting the very structure of the RIS and the randomness in the wireless channel. Moreover, we also claim that, in the context of multihop routing, the LRD-approach does not always yield the best result in terms of system performance. Finally, numerical results demonstrate the advantages of the proposed framework in terms of higher data rate, lower energy consumption, and higher energy efficiency with respect to the existing benchmark schemes. An immediate extension of this work is to investigate scenarios, where even when idle, it depends on the IUs to decide whether to act as relays to its requesting neighbours or to prioritize its own sleep mode. Moreover, we also intend to investigate the aspect and effect of imperfect channel state knowledge on the proposed strategy.

### v2

@@ III-C1 When $c=0$
This denotes a channel-aware delay unconstrained scenario, where $U_{i}$ always chooses the candidate with best channel condition, irrespective of its priority status. In case the connection does not involve any RIS, adaptive modulation (discussed in Section III-D) is employed and the corresponding data rate is determined. Note that, the scenario of considering $c=0$ and not involving any RIS, essentially represents the classical problem of channel-aware scheduling in a multiple-user scenario [39].
@@ III-F Illustrative Example for Proposed Strategy
The entire proposed strategy is presented in Fig. 2 in the form of a concise flowchart, which is illustrated below with the help of an example. As demonstrated in Fig. 1, $N$ sources $S_{1},\cdots,S_{N}$ want to communicate with their respective destination $D_{1},\cdots,D_{N}$ via IU $U_{1}$. The sources are prioritized based on their $\beta$ values, i.e., acceptable delay limit and channel condition. We consider in this work that if $U_{1}$ is idle, it agrees to act as a relay, with the higher priority users being served first. In this figure, we assume $S_{1}$ to have the highest priority and hence, it is served prior to $S_{2},\cdots,S_{N}$. Since there is no direct $U_{1}\rightarrow U_{2}$ or $U_{1}\rightarrow U_{3}$ LoS link, $U_{1}$ sends the packets to $U_{2}$ or $U_{3}$ via the RIS $R_{1,1}$. Here, we consider that the concerned RIS $R_{1}$ is already partitioned into four non-overlapping surfaces, $R_{1,1},R_{1,2},R_{1,3}$ and $R_{1,4}$, respectively. For communication purposes, $U_{1}$ selects $R_{1,1}$ to connect to $D_{1}$ via $U_{2}$ or $U_{3}$. It is interesting to note, that although LRD from $U_{3}$ to $D_{1}$ is greater than that from $U_{2}$ to $D_{1}$, $U_{3}$ is selected as the next hop. The reason for is attributed to the higher effective throughput in the case of $U_{1}\rightarrow R_{1}\rightarrow U_{3}$ link as compared to $U_{1}\rightarrow R_{1}\rightarrow U_{2}$. Finally, $U_{3}$ forwards the information packets to $U_{4}$, which in turn forwards the same to $D_{1}$ and this completes the process.
@@ IV-B Delay Bound for Complete Information Transfer for a Particular $S-D$ Pair
Finally, Remark 4 holds for the scenario, where IUs are available for each hop and the RISs in the surroundings are not being used at all. In cases where the RISs are being used due to IU unavailability, the only difference will be that the service rate $\mu_{j}$ in (35) will get replaced by $\hat{R}_{j,j+1}^{r,b}$ from (16).
@@ V-A Impact of Grouping and Spatial Correlation
By accounting for the spatial correlation, here we demonstrate the advantage of grouping in terms of the achievable data rate. A particular RIS is sub-divided into groups with identical number of patches and patch spacing being $\lambda/8$ for the correlated scenario. Moreover, only one of these groups is used for the purpose of information transfer and the results corresponding to the independent channel scenario is obtained by setting the patch spacing to $\lambda/2$. Fig. 3(a) illustrates the achievable data rate versus the number of groups into which the RIS has been sub-divided and this depicts the importance of incorporating the aspect of spatial correlation. Intuitively, this gain in performance can be understood from (III-A), i.e., grouping essentially implies combination of channels corresponding to the adjacent patches. Hence, the combination is more likely to be constructive when the channels are correlated. We also observe, that the ideal group size is somewhere in between the two extreme scenarios of considering the group size to be unity or the total number of patches in the RIS. This further corroborates the trade-off investigated in [26] between the channel estimation overhead and the power gain offered by the group.
@@
Figure 3: Impact of (a) Number of groups, (b) Patch spacing, and (c) Number of patches in a group on achievable data rate.
For a fixed number of patches in a group, Fig. 3(b) depicts the effect of the patch spacing on achievable data rate. In this context, we define the quantities $dx$ and $dy$, which denote the horizontal and vertical distance between consecutive patches, respectively. We observe that for a fixed $dx$, the achievable rate increases with decreasing $dy$ and vice-versa; for example, note the performance gap at $dx=\lambda/12$ between $dy=\lambda/12,\lambda/10,$ and $\lambda/8$. This justifies the effect of spatial correlation in (III-A), which implies that the correlation is inversely proportional to the patch spacing. On the contrary, note that this aspect of patch distance does not have any impact in the independent scenario, i.e., when $dx=dy=\lambda/2$.
@@
Fig. 3(c) shows the achievable data rate, defined in (16), as a function of the number of patches in a group. In this context, we consider groups with both optimal as well as random phases, where the optimal phase shift is calculated as in (13). Moreover, as proposed in Section III-B, we also take into account the scenario when the optimal phase shift is not a continuous quantity but it is chosen from a finite set of predetermined values. Specifically, in our proposed approach, we consider the group to be $2$ and $4$ bit coded, i.e., we choose the phase shift that is closest to the optimal from a set of $4(=2^{2})$ and $16(=2^{4})$ predetermined values, respectively. The figure demonstrates that, irrespective of the scenario, the achievable rate increases with the number of patches while the performance is worst in the case of random phases. Moreover, with an increase in the number of coded bits, the performance approaches the optimal. As can be seen, the achievable data rate with $4$ coded bits almost merges with its optimal counterpart.
@@ V-B1 Data throughput $\rm D_{T}$
Figure 5: Performance trade-off investigation.
Figure 6: Performance comparison: (a) Data throughput, (b) Energy consumption, (c) Energy efficiency.
@@ V-B3 Energy Efficiency $E_{\rm eff}$
Fig. 5 illustrates the importance of RIS grouping on the performance indicators, i.e., $\rm D_{T}$ and $\rm E_{eff}$, as a function of the number of elements in a RIS. Namely, we investigate the impact of considering a no grouping-based scenario (nGBS) and a grouping-based scenario (nGBS), where the RIS to be subdivided into four groups of equal size and the total number of patches in the RIS is varied accordingly. We observe that in both cases, $\rm D_{T}\;(\rm E_{eff})$ follows an increasing (decreasing) trend with the growing number of elements. This is because, growing number of patches supports better throughput. Note that, nGBS results in a higher data throughput as compared to GBS. This is intuitive as, unlike GBS, nGBS uses the entire RIS, which employs a fraction of the total number of patches in the RIS. Note that, nGBS limits the RIS usage to a single user, at any arbitrary point in time. On the contrary, GBS results in an enhanced energy efficiency performance. Hence, depending on the application at hand, for example, in a dense surrounding, we propose the usage of the GBS framework, which leads to better $\rm E_{eff}$ performance and also, more users can be catered to by the RISs at the same time.
@@
Fig. 6(a) depicts an overall increasing trend of $\rm D_{T}$ with number of RIS elements in a group, irrespective of the scheme considered. The reason behind this observation can be attributed to the fact that for hops involving a RIS, the achievable throughput is a function of the number of reflecting elements that are being used. As the proposed strategy is IU traffic aware, involves channel-dependent transmission techniques, and also considers the aspect of spatial correlation among the reflecting patches, it performs significantly better compared to its competitors. Moreover, the figure illustrates the fact that, as we are dealing with a delay-constrained scenario, an LRD-based approach is not always the best solution. It may result in a lesser number of hops (as observed in Fig. 4) but at the cost of reduced data rate due to factors such as wireless channel fluctuations. Furthermore, here also we observe the impact of having more number of coded bits in the RIS.
@@
Fig. 6(b) investigates the aspect of energy consumption among the proposed strategy and the existing benchmarks. It is interesting to note that ${\rm E_{C}}$, in the context of both our strategy and the LRD-based approach, decreases with an increasing number of elements in a RIS group. The reason behind this counter-intuitive observation is that, in both these scenarios, irrespective of the group size, a single phase shift is required for the entire group and the LRD becomes a criteria only when the next hop choice is concerned. Moreover, as observed from (8), the achievable data rate increases with the group size. Hence, when we combine both these factors, it leads to this interesting insight, which can also be noted from (V-B2). On the contrary, both for MRIRS and RAND, as individual phase shifts are required at all the patches, the energy consumption increases linearly with increasing group size.
@@
Finally, Fig. 6(c) shows the impact of the proposed strategy on the system energy efficiency. We observe that, while ${\rm E_{eff}}$ for MRIRS and RAND monotonically decreases with the number of reflecting elements in a group, it follows an exact opposite trend in the context of our proposed strategy. The reason behind this observation is identical to the one as already described above in the discussion related to Fig. 6(b).
As discussed earlier, specifically the work in [23] investigates the aspect of energy efficiency maximization in the context of multihop RIS-assisted wireless networks. However, we argue, that there will be significant attenuation of the transmitted signal, if only RISs and no IUs are used to connect a transmitter to its intended receiver. Hence, we propose a novel joint IU-RIS framework, which also depends on the traffic characteristics of the IUs and the associated spatial correlation at the RIS. In this context, by varying the number of reflective elements in the group, Fig. 7 compares the $\rm D_{T}$ and $\rm E_{eff}$ performance of [23] and our proposed strategy. We observe here, that irrespective of the nature of variation of these performance metrics, our proposed strategy consistently and significantly outperforms [23]. This demonstrates the importance of incorporating the idle IUs to act as relays and not solely rely on the RISs deployed in the surroundings. Moreover, this also avoids unnecessary wastage of resources, i.e., dense RIS deployment.
Figure 7: Impact of joint IU-RIS framework.
Figure 8: Impact of IUs: (a) No. of IUs vs. data throughput, (b) No. of IUs vs. energy efficiency, and (c) IU mobility vs. data throughput.
V-C Impact of IUs on System Performance
In the proposed framework, Fig. 8(a) investigates the aspect of data throughput as a function of the IU density. Here, we consider that each RIS consists of four identical groups with $\lambda/8$ spatial distance. We observe that, irrespective of the scheme, $\rm D_{T}$ increases with the growing number of IUs. This is also intuitive, as the growing number of IUs implies greater IU availability, which in turn, reduces the required number of hops for information transfer. As a result, fewer RISs are being used with increasing IU density. Moreover, it is interesting to note that, the number of RIS used in establishing the desired multihop connection asymptotically reaches zero with increasing number of IUs in the surroundings. This explains the saturating trend of $\rm D_{T}$ with the number of IUs. Furthermore, with no RISs being used with higher number of IUs, the corresponding $\rm D_{T}$ of all the considered schemes merge together, i.e., when RISs are not being used, it does not matter, whether RIS grouping is considered or not.
In an identical setup as in Fig. 8(a), Fig. 8(b) demonstrates the impact of the IU density on the system energy efficiency $\rm E_{eff}$. We observe that, for the scenarios as mentioned earlier, $\rm E_{eff}$ exhibits an increasing trend with the growing number of IUs. From (V-B2), we note that, for a device pair to communicate via a RIS, an additional power consumption is incurred due to the required phase shift operation at the RIS. On the other hand, no such power consumption takes place if two devices communicate directly. As a result, a lesser amount of energy is required, i.e., a higher $\rm E_{eff}$. Similar to the previous figure, here too, the RIS dependency reduces with the increasing number of IUs, which also explains the saturating nature of the curve.
Impact of IU Mobility
@@ V-C Impact of IUs on System Performance
Here a mobile scenario is considered, i.e., the IUs have a particular velocity in a certain direction. It is to be noted that this will have a significant effect on the performance of the proposed scheme, as the inter-IU distance changes during the process of information transfer. In this context, Fig. 8(c) illustrates the impact of IU mobility on the proposed framework, with the coverage being taken as $60$ m and the mobility aspect being modeled by the standard random waypoint (RWP) model [44]. Specifically, for this figure, we consider a pedestrian scenario [45], where we fix the maximum possible velocity $V_{\max}$ of the IUs, consider a random velocity chosen uniformly in $[0,V_{\max}]$ and observe its impact on the system data throughput. The figure illustrates that the performance deteriorates monotonically with increasing $V_{\max}$, which is intuitive. Such an observation is due to the mobility phenomenon, where a particular IU can move outside the coverage of another IU even during the communication process. This inevitably leads to outage, resulting in lesser data rate. Moreover, note that this figure with $V_{\max}=0$ is a special case corresponding to the static scenario, i.e., Fig. 6(a). Furthermore, for a fixed $V_{\max}$, ${\rm D_{T}}$ decreases with the number of coded bits, which is also inline with the observation made in Fig. 6(a) as well.
@@ VI Conclusion
In this paper, we proposed a novel priority-aware channel-dependent scheduling strategy for RIS-assisted multihop D2D communication, which also takes into account the aspect of element grouping at the RIS. The proposed strategy avoids resource wastage by exploiting the very structure of the RIS, the associated spatial correlation, and also the randomness in the wireless channel. Moreover, we also claim that, in this context of RIS-aided multihop routing, the LRD-approach does not always yield the best result in terms of system performance. Finally, numerical results demonstrate the advantages of the proposed framework in terms of higher data rate, lower energy consumption, and higher energy efficiency with respect to the existing benchmark schemes. An immediate extension of this work is to investigate scenarios, where even when idle, it depends on the IUs to decide whether to act as relays to its requesting neighbors or to prioritize its own sleep mode. Moreover, we intend to investigate the aspect of imperfect channel state and phase errors at the RISs on the proposed strategy. Finally, we also aim to comment on having various grouping criteria at the RISs and investigate its impact.